OpenStudy (horotat):

proof that any arbitrary of "C" shouldbe included all of Rational numbers ! (linear Algebra)

terenzreignz (terenzreignz):

Prove what? :/

OpenStudy (anonymous):

N<Z<Q<R<C

terenzreignz (terenzreignz):

I don't quite get it... any arbitrary <what> of "C" ?

OpenStudy (horotat):

sorry my english is not very good so i mean prove that any arbitrary subfield of "C" shouldbe included all of Rational numbers !

OpenStudy (horotat):

subfield :)

terenzreignz (terenzreignz):

Ohh... subfields... okay Isn't this more of an Abstract Algebra problem and not a Linear Algebra one? :D Nonetheless... let's proceed...

terenzreignz (terenzreignz):

@Diogo You wanna do this? :D

terenzreignz (terenzreignz):

Oh...kay... @horotat You ready to proceed?

OpenStudy (anonymous):

Well.. All imaginary numbers come from a negative square root which is an irrational number. Let's assume that \[\sqrt{2}\] is a rational number. Now lets say that there are 2 rational numbers that when divided they give you sqrt(2) \[\frac{ p }{ q } = \sqrt{2}\] \[(\frac{ p }{ q })^{2} = 2\] p^2= 2q^2 p is even so p^2 is even as well because an even number when multiplied by itself is even too. so you can say that p=2k (2k)^2 = 2q^2 4k^2 = 2q^2 2k^2 = q^2 And now you proved that q is also an even number... But that doesnt even make sense because when you devide p/q you get sqrt(2) which means sqrt(2) = irrational number. So if an imaginary number come from an irrational number and an irrational number come from rational numbers then all rational numbers are a subfield of any arbitrary C

OpenStudy (horotat):

wow, thank u all .

terenzreignz (terenzreignz):

Honestly @Diogo that made perfect sense, but it doesn't prove that any subfield of C must contain the set of Rationals... My turn :) Let F be a subfield of C. We have to show that Q is contained in F. In other words, all rational numbers would be in F. Since F is a subfield, it must contain unity (1) Since F is a subfield, it is also a group, hence the additive inverse of 1, which is -1, is also in F. Since F is a field, then any finite sum of 1's (1+1+1...+1) would also be in F. Similarly, any finite sum of -1's (-1-1-1...-1) would also be in F. This shows that all integers are in F. Since F is a field, then every nonzero element would have a multiplicative inverse, or something that you can multiply to it so that the product is unity (1). So, we have shown that all integers are in F. Then let q be a rational number. Then q = a/b, where a and b are integers. a/b = ab^-1 And since a and b^-1 are in F, so is ab^-1 which is equal to a/b which is equal to q. Thus, for any rational number q, q would be in F. Thus, Q (the set of Rational numbers) would be contained in F. Q(uite) E(asily) D(one) LOL I hope you guys enjoyed that :)

OpenStudy (horotat):

@Diogo thanks alot :) & @terenzreignz THANK you more alot :D openstudy is really relly great for me ...

terenzreignz (terenzreignz):

Stick around, @horotat and see if you can be of help to others yourself :) Welcome to OpenStudy, and enjoy :D

OpenStudy (horotat):

:) thanks again, i hope that i'll be helpful